OpenStudy (anonymous):

The pair of points is on the graph of an inverse variation. Find the missing value. (9, 5) and (x, 6)

OpenStudy (cwrw238):

5 = k/9 (inverse variation y = k/x k = 45 6 = 45/x x = 45/6 = 7.5